From Earth we can measure the radius of Mars using our telescopes. An estimate for it is 3.39 x 106 m. By sending an exploratory robot to Mars, we determined the acceleration due to gravity on its surface as 3.73 m/s2. Estimate the mass of Mars
1
Expert's answer
2020-12-10T10:59:09-0500

"F = m_1*a"

"\text{where }m_1 \text{ -arbitrary body mass }a =3.73 \text{acceleration due to gravity}"

"\text{on the other hand, according to the law of universal gravitation:}"

"F = G\frac{m_1M}{R^2}"

"\text{where } M\text{ planet mass}"

"R =3.39*10^6\text{ distance between bodies}"

"G \approx6.67*10^{-11}\text{ gravitational constant}"

"m_1*a = G\frac{m_1M}{R^2}"

"M= \frac{a*R^2}{G}= \frac{3.73*3.39^2*10^{12}}{6.67*10^{-11}}\approx6.42*10^{23}kg"

Answer: "M \approx6.42*10^{23}kg"